undeadknight26 (undeadknight26):

Which of the following describes the intercepts on the graph? The initial acceleration of the runner was 8 m/s2 and the runner stopped after 6 seconds. The initial acceleration of the runner was 6 m/s2 and the runner stopped after 8 seconds. The initial velocity of the runner was 8 m/s and the runner stopped after 6 seconds. The initial velocity of the runner was 6 m/s and the runner stopped after 8 seconds.

OpenStudy (anonymous):

The graph?

undeadknight26 (undeadknight26):

wrong question sorry.

OpenStudy (solomonzelman):

The intersect the axis. y-intercept intersects the y-axis x-intercept intersects the x-axis

undeadknight26 (undeadknight26):

real question. Which of the following best defines 4 to the power of 2 over 3? Square root of 16 Cube root of 4 Square root of 4 Cube root of 16

OpenStudy (anonymous):

4^2 = 16/3 = 5.3 sq rt of 16 = 4 cube root of 4 = 64 sq rt of 4 = 2 cube root of 16 = 4096